Accelerator Update for January 28, 2009 - January 30, 2009 |
Wednesday Jan 28 The day shift began with Operations monitoring the Tevatron (TeV) store 6757 (197.91E30) and the Recycler stash (112.7E10), with the Antiproton Source stacking (16.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE experiment taking beam. NuMI was down for an access to MI65. At 12:56 PM I:KPS1A (MI-10 proton injection kicker) tripped off. Experts were called to investigate. Beam was off I:KPS1A was back on at 4:00 PM. Experts had to drain and repair a fluorinert leak on the thyratron for I:KPS1A. Beam remained off due to studies on LRF4 and an access into the AP0 water cage. Beam returned to Pbar, NuMI and MiniBooNE at 5:07 PM. Thursday Jan 29 The midnight shift began with Operations monitoring store 6757 (61.41E30) and the stash (362.6E10), with Pbar stacking (35.0E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6757 at 12:47 AM. Operations put the TeV into shot setup at 1:00 AM. Operations established store 6758 at 2:25 AM with an initial luminosity of 330.7E30. At 10:02 AM, NuMI was off while experts replaced a T0 card inside E:V100. NuMI beam returned at 10:51 AM. Linac RF Station #3 (LRF3) went down at 5:19 PM. Experts investigated, no beam for stacking or to the users. Experts replaced a bad transformer for LRF3, beam returned to users at 7:01 PM. Operations terminated store 6758 at 7:35 PM. D0 began a controlled access at 8:13 PM. The Main Injector had a watchdog trip at 9:56 PM. After rebooting the Main Injector Low Level RF and damper systems, beam returned at 10:55 PM. D0 came out of access at 10:55 PM. CDF had requested a controlled access and began the access at 11:37 PM. Friday Jan 30 The midnight shift began with the CDF in access, with Operations monitoring the Recycler stash (380.1E10), with the Antiproton Source stacking (98.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the D-Zero (D0) experiment waiting for the next store, and with the MiniBooNE and NuMI experiments taking beam. CDF came out of access at 3:00 AM. Operations put the TeV into shot setup at 5:23 AM. Operations established store 6762 at 6:37 AM with an initial luminosity of 340.17E30. The Plan for today and the Weekend. The plans are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 01/30/2009 email Fermilab |
| Security, Privacy, Legal | |